Marin Products produces three products — DBB-1, DBB-2, and DBB-3 from a joint process. Each product may be sold at the split-off point or processed further. Additional processing requires no special facilities, and production costs of further processing are entirely variable and traceable to the products involved. Key information about Marin's production, sales, and costs follows.

DBB-1 DBB-2 DBB-3 Total
Units Sold 14,000 22,000 30,000 66,000
Price (after addt’l processing) $ 30 $ 15 $ 40
Separable Processing cost $ 441,000 $ 180,000 $ 263,000 $ 884,000
Units Produced 14,000 22,000 30,000 66,000
Total Joint Cost $ 4,100,000
Sales Price at Split-off $ 5 $ 15 $ 35

The amount of joint costs allocated to product DBB-1 using the physical measure method is (calculate all ratios and percentages to 2 decimal places, for example 33.33%, and round all dollar amounts to the nearest whole dollar):

Multiple Choice

  • $16,275.

  • $869,610.

  • $1,863,450.

  • $14,760.

  • $1,366,530.
Answer #1

Percentage = 14000/66000 = 21.21%

Joint cost allocated to product DBB-1 = Joint cost* percentage share

= 4100000*21.21%

Joint cost allocated to product DBB-1 = 869610

So answer is b) $869610